# LSR-MCTS: Alleviating Long Range Dependency in Code Generation

### v3

@@ 5.1 Experimental Setup
Dataset Three commonly public Python code datasets are chosen for comparative analysis, including HumanEval [^2] https://huggingface.co/datasets/openai/openai_humaneval Chen et al. (2021) and MBPP [^3] https://huggingface.co/datasets/google-research-datasets/mbpp Austin et al. (2021) of foundational difficulty and Code Contests [^4] https://huggingface.co/datasets/deepmind/code_contests Li et al. (2022) of competitive programming difficulty. Each dataset entry comprises a natural language description of a programming problem, associated test cases, and manually crafted solutions. The HumanEval and MBPP datasets boast sizes of 164 and 500, respectively, yet they lack a clear distinction between public and private test cases. To solve this problem, the test cases for each data are evenly divided into two portions: one serving as public test cases for the evaluation and refinement during LSR-MCTS; the other as private test cases for the assessment of experimental performance. The Code Contests dataset has 165 programming competition problems curated from Codeforces Mishra et al. (2021) and CodeNet Puri et al. (2021). Each problem is accompanied by clearly defined public and private test cases, indicating that no further data processing is required.
Models Two categories of LLMs are utilized to evaluate our proposed method. The first category is public code-specific LLMs, including CodeLlama Roziere et al. (2023) and aiXcoder, which are state-of-the-art in Code LLMs. We employ aiXcoder and the instruction fine-tuned version of CodeLlama, both with 7 billion parameters. To demonstrate the generalizability of LSR-MCTS, extensive experiments are also conducted on general LLMs. We choose GPT-4 Achiam et al. (2023), one of the most well-known models. Additionally, the recently trained Llama3 Dubey et al. (2024), which is shown to have significant advantages across various tasks, is also included in the experiments. The 8 billion parameters fine-tuned version is selected.
Baselines The baselines are categorized into three groups: traditional decoding methods, self-refine methods, and MCTS-based methods. Beam-search and top-p are chosen as the traditional baselines, which are widely used in generation tasks. For the self-refine method, Reflexion Shinn et al. (2024) is adopted. Reflexion continuously refines and regenerates the code based on the textual feedback from LLMs until the evaluation metrics reach a plateau. For the MCTS-based method, PG-TD Zhang et al. (2023) is selected, which is a token-level MCTS approach. To compare PG-TD and LSR-MCTS, the hyperparameter c in Equation 1 is set to 4, and the rollout n in Algorithm 1 is set to 100.
Methods HumanEval MBPP Code Contests
pass@1 pass@3 pass@5 pass@1 pass@3 pass@5 pass@1 pass@3 pass@5
CodeLlama-7B-Instruct
Beam-Search 36.1 39.4 40.2 30.1 32.9 33.6 6.7 8.8 9.5
T-MCTS 42.2 46.3 47.9 36.6 38.3 40.7 8.9 10.0 11.1
TSR-MCTS 43.5 47.4 48.2 37.8 39.0 41.6 9.3 10.7 11.5
L-MCTS 43.8 48.1 48.3 38.5 40.7 42.5 9.2 10.4 11.1
LSR-MCTS 45.7 49.4 50.6 40.8 42.2 43.9 9.8 11.2 12.0
GPT-4
Beam-Search 85.4 86.7 87.2 49.1 49.9 50.2 12.3 14.3 15.2
T-MCTS 89.3 90.7 91.3 53.2 54.4 55.8 14.7 15.4 16.3
TSR-MCTS 89.7 90.9 91.5 53.6 54.8 55.9 14.9 15.9 16.8
L-MCTS 89.7 91.4 92.4 53.7 54.6 56.3 15.2 16.0 16.8
LSR-MCTS 90.6 92.3 93.1 54.9 55.8 57.1 16.2 17.3 17.9
Table 2: Ablation performance of different components in LSR-MCTS.
Metric The unbiased estimation of $pass@k$ Chen et al. (2021) is used to assess the functional correctness of code generated by LLMs, where $k$ code samples are produced for each problem, with $k=1,3,5$ serving as the evaluation criteria. We generate $n\geq k$ programs for each data, and the number of programs $c$ that pass the private test cases is calculated, thereby determining the unbiased estimate:
$pass@k:=\underset{Problems}{\mathbb{E}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right]$ (3)
@@ 5.2 Main Experiments
A more in-depth analysis of the dataset reveals that these models demonstrate greater enhancement on the challenging competitive programming dataset Code Contests, as opposed to the normal difficulty of HumanEval and MBPP. Particularly noteworthy is the significant improvement of 12.4% for aiXcoder-7B at pass@5, indicating that LSR-MCTS is adept at stimulating the potential of LLMs on complex issues.
As the value of $k$ in $pass@k$ changes, the proportion of enhancement by LSR-MCTS for the same model and dataset remains generally stable. The unbiased estimation characteristic of $pass@k$ suggests that the model exhibits excellent robustness.
5.3 Ablation Results
We designed experiments to analyze the influence of the two integral components of LSR-MCTS on model performance. Table 2 shows a comparative analysis, where T-MCTS eliminates the line-level strategy and only considers token-level MCTS. TSR-MCTS adds a self-refine mechanism on the basis of token-level. In contrast, L-MCTS removes the self-refine module from LSR-MCTS. Here, the MCTS rollout is set to 100 for all.
The comparison results reveal that, under identical hyperparameter settings, both components are instrumental in enhancing the performance of the decoding strategy, proving indispensable in their respective roles. Interestingly, the contribution of the self-refine module at the token-level is negligible, with almost no discernible improvement. This stands in sharp contrast to the significant enhancement observed at the line-level, probably because the token-level approach may diminish the semantic connections between tokens over long distances, which is precisely what self-refine needs to exploit. Therefore, the combination of the two does not yield benefits. This observation emphasizes the synergy between line-level MCTS and the self-refine mechanism, which can reinforce each other effectively.
Appendix B presents experiments with different self-refine methods, demonstrating that the self-refine mechanism can effectively harness the capabilities of LLMs.
@@ 5.4 Parameter Analysis
Figure 4: The impact of hyperparameter variations on GPT-4 performance. Hyperparameters include the number of max rollouts $n$, the UCT parameter $c$, and the maximum number of child nodes $m$ in the tree.
@@ 6 Conclusion
In this paper, we present a novel non-training decoding strategy called LSR-MCTS, which leverages the characteristics of inter-token attention to prove that line-level units are more effective for code generation. This strategy consists of a line-level MCTS and a self-refine mechanism. The former reconstructs the content of each node, segmenting the entire code block into context, line, and supplement, finalizing the line incrementally from a global perspective as the depth of the node increases. To mitigate the impact of potential branch defects caused by the limited number of child nodes, the self-refine mechanism is employed to discover more effective programs and rectify code blocks. Extensive experiments conducted on three public code generation datasets demonstrate that LSR-MCTS achieves state-of-the-art performance across all models.
Limitations
In this section, we discuss two limitations of LSR-MCTS. On the one hand, although using lines as the basic processing unit can reduce the number of operations during decoding, the self-refine mechanism calls LLMs at each node, making the time advantage negligible. On the other hand, code generation in the real world typically involves requirements without test cases. The dependency on public test cases needs to be mitigated by leveraging existing datasets or employing LLMs to generate them automatically. Therefore, improving time efficiency and tackling the scarcity of test cases are two urgent challenges that demand attention.
Appendix A Prompts in LSR-MCTS
The prompts for LLMs policy in LSR-MCTS are shown below.
A.1 The generation prompt PR
1Complete the Python function below by adding the necessary code to achieve the description in the functions docstring. Please return the whole function code without any description.
2
3### Programming Problem:
4{NL Description}
5
6### Solution:
A.2 Self-refine prompt SRP
1As a Python writing assistant, you will receive a natural language problem description, a complete function implementation generated by LLMs, a series of public test cases, and the line attribute of the current node (denoted as L). Please verify whether the function implementation is correct and if it can pass all the public test cases. If it is correct, return the function implementation; otherwise, retain the code before L and modify the code snippet after L (including L). The self-refine input follows the format below, please return the entire function code without any description:
2
3### Programming Problem:
4{NL Description}
5
6### Function Implementation:
7{Generated Code}
8
9### Public Test Cases:
10{Public Test Cases}
11
12### Node Line:
13{Line}
14
15### Refined code:
A.3 Feedback prompt FP
1As a programming scoring assistant, you will receive a natural language problem description, a complete function implementation generated by LLMs, and a series of public test cases. Please score the program based on the simplicity of the function, the correctness of its functionality, and its ability to pass public test cases. The returned value should be an integer from 1 to 10, format: "[[score]]", for example: "Feedback Score: [[5]]". The input data format you received is as follows:
2
3### Programming Problem:
4{NL Description}
5
6### Function Implementation:
7{Generated Code}
8
9### Public Test Cases:
10{Public Test Cases}
11
12### Feedback Score:
Appendix B Analysis of Different Self-Refine Methods
To substantiate the superiority of self-refine and to compare different refine approaches, in addition to the node expansion method mentioned in Section 4.2 and Algorithm 1 Line 13-15, we also introduce a new one based on feedback scoring, called LFS-MCTS. It replaces Line 13 of Algorithm 1, utilizing novel scoring feedback as supplementary information for each node. To achieve this, we designed a feedback prompt FP, which is shown in Appendix LABEL:lstlisting:FP. Utilizing FP to input into LLMs to obtain $score=M_{\theta}(FP+next\_node,1)$, which is then integrated into the reward function for subsequent evaluation and the backpropagation phase:
$values=\lambda score+R(next\_node)$ (4)
where $\lambda$ is the weight for the feedback score.
Figure 5: Performance of various self-refine methods on CodeLlama-7B-Instruct and GPT-4.
Analysis of the data from Figure 5 indicates that both self-refine methods positively impact the performance of line-level MCTS. The feedback scoring method shows significant room for improvement in CodeLlama, whereas the enhancement is minimal in GPT-4. This suggests that when the original method already has a high accuracy rate, the feedback scoring may not substantially increase the search space as effectively as LSR-MCTS does, to find the correct solutions. Comparative experiments confirm the efficiency of the LSR-MCTS approach.

### v4

@@ 4.1 Experimental Setup
Dataset. Three commonly public Python code datasets are chosen for analysis, including HumanEval [^1] https://huggingface.co/datasets/openai/openai_humaneval [5] and MBPP [^2] https://huggingface.co/datasets/google-research-datasets/mbpp [3] of foundational difficulty and Code Contests [^3] https://huggingface.co/datasets/deepmind/code_contests [19] of competitive programming difficulty. Each dataset comprises a natural language description of a programming problem, associated test cases and manually crafted solutions.
Models. Two categories of LLMs are utilized to evaluate our proposed method. The first category is public code-specific LLMs, including CodeLlama-7B-Instruct [24] and aiXcoder-7B. To demonstrate the generalizability of LSR-MCTS, extensive experiments are also conducted on general LLMs GPT-4 [1] and Llama3-8B-Instruct [9].
Baselines are categorized into three groups: traditional decoding methods, self-refine methods, and MCTS-based methods. Beam-search and Top-p are chosen as the traditional baselines, which are widely used in generation tasks. For the self-refine method, Reflexion [25] is adopted. It continuously refines and regenerates the code based on the textual feedback from LLMs until convergence. For the MCTS-based method, PG-TD [34] is selected, which is a token-level MCTS approach. The hyperparameter c in Equation 1 is set to 4, and the rollout n is set to 100 for comparison.
Metric. $pass@k$ [5] is used to assess the functional correctness of code generated by LLMs, where $k$ code samples are produced for each problem, with $k=1,3,5$. We generate $n\geq k$ programs for each data, and $c$ of them pass the private test cases. The unbiased estimate is calculated:
$pass@k:=\underset{Problems}{\mathbb{E}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]
@@ 4.2 Main Experiments
A more in-depth analysis of the dataset reveals that these models demonstrate greater enhancement on the challenging competitive programming dataset Code Contests, as opposed to the normal difficulty of HumanEval and MBPP. Particularly noteworthy is the significant improvement of 12.4% for aiXcoder at pass@5, indicating that LSR-MCTS is adept at stimulating the potential of LLMs on complex issues.
As the value of $k$ changes, the proportion of enhancement by LSR-MCTS for the same model and dataset remains generally stable. The unbiased estimation characteristic of $pass@k$ suggests that the model exhibits excellent robustness.
4.3 Ablation Study
We designed experiments to analyze the influence of the two integral components of LSR-MCTS on model performance. Table 1 shows a comparative analysis on CodeLlama and GPT-4. T-MCTS eliminates the line-level strategy and only considers token-level MCTS, i.e., PG-TD. TSR-MCTS adds a self-refine mechanism on the basis of token-level. In contrast, L-MCTS removes the self-refine module from LSR-MCTS. Here, the MCTS rollout is set to 100 for all.
The comparison results reveal that, under identical hyperparameter settings, both components are instrumental in enhancing the performance of the decoding strategy. Interestingly, the contribution of the self-refine module at the token-level is negligible, with almost no discernible improvement. This stands in sharp contrast to the significant enhancement observed at the line-level, probably because the token-level approach may diminish the semantic connections between tokens over long distances, which is precisely what self-refine needs to exploit. Therefore, the combination of the two does not yield benefits. This observation emphasizes the synergy between line-level MCTS and the self-refine mechanism, which can reinforce each other effectively.
@@ 4.4 Parameter Analysis
Fig. 4: The impact of hyperparameter variations on GPT-4 performance. Hyperparameters include the number of max rollouts $n$, the UCT parameter $c$, and the maximum number of child nodes $m$ in the tree.
@@ 5 Conclusion
In this paper, we present a novel training-free decoding strategy called LSR-MCTS. This strategy consists of a line-level MCTS and a self-refine mechanism. The former segments the code block of each node into context, line, and supplement, generating the code line-by-line from a global perspective. The self-refine mechanism is employed to discover more effective programs and rectify code blocks. Extensive experiments conducted on three benchmarks demonstrate that LSR-MCTS achieves state-of-the-art performance across all models.